OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17 If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17 @Mathematics

OpenStudy (anonymous):

(–2y + 2) + 3y = 17 x + 3(2y + 2) = 17 (2y + 2) + 3y = 17 x + 3(–2y + 2) = 17

OpenStudy (eyust707):

x= 2 + 2y

OpenStudy (anonymous):

x - 2y = 2 x = 2 + 2y 2 + 2y + 3y = 17 2 + 5y = 17 2 = 12y y = 1/6